OpenStudy (anonymous):

There is an inverse relationship between x and y. If y is twenty when x is 2, what is x when y is 8? 5 8 40 80

OpenStudy (anonymous):

\[y = k/x\] \[20 = k /2\] \[40 = k\] therefore \[y = 40/x\] \[8 = 40/x\] \[x = 40/8\] \[x = 5\]

OpenStudy (anonymous):

@alexandercpark ty